AgriGrow is to purchase a tractor for over-the-road hauling for $90,000. It is expected to be of use to the company for 6 yeaUse double declining balance depreciation (no half-year convention, no switching). End of Year ATCF -90000 34200 30200 27533.
Answer #1

I have solved it using excel refer the picture below

fx =MIRR( U10:U14 ,10%,10%) V W 10 11 12 13 EoY ATCE 0 -90000 1 34200 30200 27533.33 36466.67 After Tax PW | $11,643.01 After

Manually we can derive it as follows

First of all convert the cash outflow into present value

Present value of outflow = $ 90,000

Now convert inflow into future value

FV = 34,200*1.1^3 + 30,200*1.1^2 + 27,533.33*1.1 + 36,466.67 = $ 148,815.63

Now equate these two values in the following manner

90,000(1+ ERR)4 = 148,815.63

(1 + ERR)4 = 148,815.63 90.000

(1 + ERR)* = 1.653505

ERR = 1.6535057 – 1

ERR = 1.133969 - 1

ERR = 0.133969

ERR = 13.397%